OpenStudy (anonymous):

Please help me with algebra 2 homework? Another measure of variation is variance, which equals O2 .Find the variance and standard deviation of the data set. Round to the nearest tenth. 7, 8, 4, 10, 6, 10 Select one: a. variance = 4.6; standard deviation = 2.3 b. variance = 1.5; standard deviation = 2.1 c. variance = 1.5; standard deviation = 2.3 d. variance = 4.6; standard deviation = 2.1

OpenStudy (loser66):

the last one. I give you my trick. take \(\sqrt{variance}\) it is standard deviation \(\sqrt{4.6}=2.1\) so, the last one is the answer Let's check others: b) \(\sqrt{1.5}= 1.2\) , your option b is 1.5 and 2.1, not good c) the same,